# Domain Generalization and Adaptation in Intensive Care with Anchor Regression

### v1

@@ 4.3 Selecting $\gamma$ is difficult
In figures˜6, 5, 4, 8 and 7 and figures˜12, 13, 14, 15, 16, 17, 18, 19 and 20 in appendix˜C, we mark the value of $\gamma$, selected for the default choice of hyperparameters by leave-one-environment-out cross-validation (LOEO-CV), minimizing the average OOD MSE for regression and the average OOD negative log-likelihood for classification, with a star.
@@
As anchor regression is meant to improve worst-case, not average-case performance to new environments, one could argue to minimize the worst-case (instead of average) MSE or negative log-likelihood over the OOD targets with LOEO-CV. However, as discussed in section˜4.2, varying noise levels between domains make a comparison between them difficult.
Finally, Rothenhäusler et al. (2021) suggest, as an alternative to LOEO-CV, to select $\gamma$ based on prior knowledge about shift sizes. As we see in section˜4.4, such an approach can be problematic for linear anchor regression, where we observe an interaction between conventional regularization and the optimal value for $\gamma$. We discuss this in the next section.
@@ 4.4 On the choice of other hyperparameters
As mentioned in section˜3.1, we add an elastic net regularization term of the form $\lambda\left(\eta\|\beta\|_{1}+(1-\eta)\|\beta\|_{2}^{2}\right)$ to equations˜1 and 2 for linear anchor regression. Let $\lambda_{\mathrm{max}}$ be minimal such that all parameters other than the intercept of a lasso model ($\eta=\gamma=1$) with $\lambda=\lambda_{\mathrm{max}}$ are zero. In figure˜4, we show the performance of anchor regression predicting log(creatinine) in 24 hours with $\eta=0,0.5,1$ and $\lambda=10^{-2}\lambda_{\mathrm{max}},10^{-3}\lambda_{\mathrm{max}},10^{-4}\lambda_{\mathrm{max}}$. We observe an interaction between $\gamma$ and $\lambda$: As we increase the amount of conventional regularization $\lambda$ by a factor of 10, the optimal amount of anchor regularization $\gamma$ decreases by around 10. We observe the same effect when predicting log(lactate) in 4 hours, see figure˜18 in appendix˜C.
@@
Due to the algorithm’s increased variance, the corresponding plots for nonlinear anchor boosting are more difficult to interpret. In figure˜6, we show anchor boosting’s AuPRC predicting circulatory failure within 8 hours. We vary the number of trees from 500, 1000, to 2000, and the trees’ maximal depth from 2, 3, to 4. We do not observe a clear effect that a lower value of $\gamma$ being more optimal for a more strongly conventionally regularized boosted model. In contrast, for a fixed maximal depth, the shape of the models’ AuPRC with a varying number of trees stays similar. This suggests that anchor regularization has mainly an effect in the early boosting iterations. We observe the same behavior when predicting log(lactate), log(creatinine), and acute kidney injury, see figures˜13 and 12 in appendix˜C.
@@ 4.5 Which variables are good choices to use as anchors?
Results are similar when predicting log(creatinine), acute kidney injury, and circulatory failure, see figures˜15, 16, 17 and 14 in appendix˜C.
@@ 4.6 Refitting models using few samples from the target
We apply the refitting methodology of sections˜3.4 and 3.5 to the ICU data. For 20 seeds, we draw increasing subsets of sizes 25, 35, 50, 70, 100, … of patient IDs from the target’s train set and then refit the anchor models on these patients. We select $\gamma\in\{1,2,4,\ldots\}$, the decay rate $\mathrm{dr}\in\{0,0.2,\ldots,1\}$, and the prior’s width $\alpha\in\{10,\sqrt{10},\ldots,10^{-3}\}$ with 5-fold cross-validation on the sampled target data. We also fit models using only the sampled target data. For each seed, we fit monotonous cubic splines to the model performances and use these to extrapolate beyond the number of patient samples available for each domain. We show the performance of these approaches predicting acute kidney injury within 48 hours in figure˜9 and predicting log(creatinine), log(lactate), and circulatory failure in figures˜21, 22, 23, 24, 25, 26 and 27 in appendix˜C.
@@ 4.7 Three regimes and information in external data
We described the three regimes and their transition points in section˜3.6. We summarize the empirical transition points for anchor boosting applied to acute kidney injury, circulatory failure, log(creatinine), and log(lactate) prediction in figure˜10. We believe that these results are important to understand the potential of using external data for domain adaptation. We show the equivalent plot for linear anchor and logistic regression in figure˜28 in appendix˜C.
@@
Figure 10: Regime transitions for boosted tree models as described in section˜3.6 and figure˜2.  denotes the regime transition i $\to$ ii,   the regime transition ii $\to$ iii, and  denotes the external data’s value.
5 Conclusion
@@
We also propose a framework that quantifies the value of external out-of-distribution data. The framework quantifies the value of external data by determining the number of target-domain samples required to match an OOD model’s performance. It uses this to identify three distinct regimes of data utility: domain generalization, domain adaptation, and target-only training. This taxonomy provides a data-driven methodology for practitioners to decide how to integrate external data. Our results show that while external data is valuable when target data is scarce, its value diminishes and it eventually becomes obsolete the more in-distribution data becomes available.
Some challenges remain. The selection of the anchor regularization parameter $\gamma$ is non-trivial and interacts with conventional regularization. Also, the choice of the anchor variable remains subjective and a more data-driven methodology is missing. Finally, our proposed framework for evaluating external training data utility is highly general and we would welcome its application to other (multi-source) domain adaptation learning problems.
@@
Our study uses harmonized data and resources expanding the ricu R-package (Bennett et al., 2023) from prior work by Burger et al. (2024, 2025), who describe a large-scale multi-center ICU dataset. This extended version of ricu includes additional code and configuration files to harmonize data from additional data sources (NWICU, PICdb, and Zigong) and to extract additional variables and will be published together with Burger et al. (2025). See https://github.com/ratschlab/icarefm for details.
@@ Acknowledgements
Malte Londschien is partially supported by the ETH Foundations of Data Science. We gratefully acknowledge early access to the harmonized dataset and ricu resources provided by Burger et al. (2024, 2025). We would like to thank Cyrill Scheidegger, Dinara Veshchezerova, Dominik Rothenhäusler, Gianna Wolfisberg, Hugo Yèche, Jonas Peters, Julia Kostin, Markus Ulmer, Maybritt Schillinger, Michael Law, Nicolas Bennet, Niklas Pfister, Olga Mineeva, Patrick Rockenschaub, Paola Malsot, Sorawit Saengkyongam, and Yuansi Chen for helpful discussions and comments.
@@ A.1 Tasks
Acute kidney injury is defined as AKI 3 according to the KDIGO guidelines (Acute Kidney Injury Work Group, 2012). These are: (i) A patient has an increase in creatinine of a factor of 3 relative to their 7 day baseline; (ii) A patient has acute kidney injury level 1 according to the KDIGO guidelines and has a creatinine value of at least 4.0 mmol/L; (iii) A patient has anuria, that is, has not produced urine over the last 12 hours; (iv) A patient had an average relative urine rate below 0.3ml/kg/h over the last 24 hours; (v) The patient has started renal replacement therapy. We say that a patient is experiencing an acute kidney injury event if at least one of the conditions (i) - (v) apply. As urine rate is only measured sparsely for NWICU, PICdb, and NWICU, we define a negative kidney injury event as (i, ii) creatinine is measured and low, (iii, iv) urine rate is measured and normal, or not measured, (v) the patient is not receiving renal replacement therapy. Again, requiring that urine rate is measured and normal would result in very few negative events for the datasets where urine rate is only very sparsely measured.
@@ A.2 Feature engineering
We use an 8-hour horizon for feature engineering for the prediction tasks of log(lactate) and circulatory failure and a 24-hour horizon for the feature engineering for the prediction tasks of log(creatinine), acute kidney injury.
@@ B.1 Details on the probit anchor loss
Recall that $\Phi$ and $\varphi$ are the Gaussian distribution’s cumulative distribution function and probability density function. For scores $f\in\mathbb{R}^{n}$ and outcome $y\in\{-1,1\}^{n}$, the negative log-likelihood is $-\sum_{i=1}^{n}\log(\Phi(y_{i}f_{i}))$ with gradient $r(f):=-y\cdot\varphi(f)/\Phi(yf)$. We use the gradient as (score) residuals (Kook et al., 2022). The probit anchor loss with parameter $\gamma$ is
@@
[…]
@@ B.2 The logistic anchor loss is not convex
Write $\sigma(f)=1/(1+e^{-f})$. For scores $f\in\mathbb{R}^{n}$ and outcomes $y\in\{-1,1\}^{n}$, the negative log-likelihood using the logistic link $\sigma$ is $-\sum_{i=1}^{n}\log(\sigma(y_{i}f_{i}))$ with gradient $r(f):=y\cdot\sigma(-yf)$. Using the gradient as (score) residuals (Kook et al., 2022), the logistic anchor loss with parameter $\gamma$ is
@@ Appendix C Additional figures
Figure 28: Regime transitions for linear models as described in section˜3.6 and figure˜2.  denotes the regime transition i $\to$ ii,   the regime transition ii $\to$ iii, and  denotes the external data’s value.

### v2

@@ 4.3 Selecting $\gamma$ is difficult
In figures˜6, 5, 4, 8 and 7 and figures˜12, 13, 14, 15, 16, 17, 18, 19 and 20 in appendix C, we mark the value of $\gamma$, selected for the default choice of hyperparameters by leave-one-environment-out cross-validation (LOEO-CV), minimizing the average OOD MSE for regression and the average OOD negative log-likelihood for classification, with a star. We prefer the negative log-likelihood over AuPRC for model selection. As a proper scoring rule, the log-likelihood is sensitive to calibration and thus better captures overfitting than rank-based metrics (Friedman et al., 2000).
@@
Anchor regression aims to improve worst-case, not average-case performance to new environments. Thus, one could argue to minimize the worst-case (instead of average) MSE or negative log-likelihood over the OOD targets with LOEO-CV. However, as discussed in section˜4.2, varying noise levels between domains make a comparison between them difficult.
Finally, Rothenhäusler et al. (2021) suggest to select linear anchor regression’s $\gamma$ based on prior knowledge about shift sizes, as an alternative to LOEO-CV. In section˜4.4, we observe an interaction between conventional regularization and the optimal value for $\gamma$ of linear anchor regression, making such an approach problematic.
The interpretation of $\gamma$ as the relative shift magnitude in the test domain is theoretically grounded for linear anchor regression. While we intuitively expect a similar relationship in the nonlinear case, there are no formal guarantees. Empirically, $\gamma$ serves as a parameter to achieve best worst-case out-of-distribution performance, making LOEO-CV a practical tool for model selection.
@@ 4.4 On the choice of other hyperparameters
As mentioned in section˜3.1, we add an elastic net regularization term of the form $\lambda\left(\eta\|\beta\|_{1}+(1-\eta)\|\beta\|_{2}^{2}\right)$ to equations˜1 and 2 for linear anchor regression. Let $\lambda_{\mathrm{max}}$ be minimal such that all parameters other than the intercept of a lasso model ($\eta=\gamma=1$) with $\lambda=\lambda_{\mathrm{max}}$ are zero. In figure˜4, we show the performance of anchor regression predicting log(creatinine) in 24 hours with $\eta=0,0.5,1$ and $\lambda=10^{-2}\lambda_{\mathrm{max}},10^{-3}\lambda_{\mathrm{max}},10^{-4}\lambda_{\mathrm{max}}$. We observe an interaction between $\gamma$ and $\lambda$: As we increase the amount of conventional regularization $\lambda$ by a factor of 10, the optimal amount of anchor regularization $\gamma$ decreases by around 10. We observe the same effect when predicting log(lactate) in 4 hours, see figure˜18 in appendix C.
@@
Due to the algorithm’s increased variance, the corresponding plots for nonlinear anchor boosting are more difficult to interpret. In figure˜6, we show anchor boosting’s AuPRC predicting circulatory failure within 8 hours. We vary the number of trees from 500, 1000, to 2000, and the trees’ maximal depth from 2, 3, to 4. We do not observe a clear effect that a lower value of $\gamma$ being more optimal for a more strongly conventionally regularized boosted model. In contrast, for a fixed maximal depth, the shape of the models’ AuPRC with a varying number of trees stays similar. This suggests that anchor regularization has mainly an effect in the early boosting iterations. We observe the same behavior when predicting log(lactate), log(creatinine), and acute kidney injury, see figures˜13 and 12 in appendix C.
@@ 4.5 Which variables are good choices to use as anchors?
Results are similar when predicting log(creatinine), acute kidney injury, and circulatory failure, see figures˜15, 16, 17 and 14 in appendix C.
@@ 4.6 Refitting models using few samples from the target
We apply the refitting methodology of sections˜3.4 and 3.5 to the ICU data. For 20 seeds, we draw increasing subsets of sizes 25, 35, 50, 70, 100, … of patient IDs from the target’s train set and then refit the anchor models on these patients. We select $\gamma\in\{1,2,4,\ldots\}$, the decay rate $\mathrm{dr}\in\{0,0.2,\ldots,1\}$, and the prior’s width $\alpha\in\{10,\sqrt{10},\ldots,10^{-3}\}$ with 5-fold cross-validation on the sampled target data. We also fit models using only the sampled target data. For each seed, we fit monotonous cubic splines to the model performances and use these to extrapolate beyond the number of patient samples available for each domain. We show the performance of these approaches predicting acute kidney injury within 48 hours in figure˜9 and predicting log(creatinine), log(lactate), and circulatory failure in figures˜21, 22, 23, 24, 25, 26 and 27 in appendix C.
@@ 4.7 Three regimes and information in external data
We described the three regimes and their transition points in section˜3.6. We summarize the empirical transition points for anchor boosting applied to acute kidney injury, circulatory failure, log(creatinine), and log(lactate) prediction in figure˜10. We believe that these results are important to understand the potential of using external data for domain adaptation. We show the equivalent plot for linear anchor and logistic regression in figure˜28 in appendix C.
@@
Figure 10: Regime transitions for boosted tree models as described in section˜3.6 and figure˜2. denotes the regime transition i $\to$ $i$,   the regime transition $i$ $\to$ iii, and denotes the external data’s value.
5 Discussion
@@
We also propose a framework that quantifies the value of external out-of-distribution data. The framework quantifies the value of external data by determining the number of target-domain samples required to match an OOD model’s performance. It uses this to identify three distinct regimes of data utility: domain generalization, domain adaptation, and target-only training. This taxonomy provides a data-driven methodology for practitioners to decide how to integrate external data. Our results show that while external data is valuable when target data is scarce, its value diminishes and it eventually becomes obsolete as more in-distribution data becomes available.
Some challenges remain. The empirical results are not uniformly positive across all settings. Furthermore, the selection of the anchor regularization parameter $\gamma$ is non-trivial and interacts with conventional regularization. Finally, the choice of the anchor variable remains subjective and a more data-driven methodology is missing.
However, given the inherent difficulty of domain generalization in the multi-center ICU setting, the proposed framework offers a path forward. Using a distributionally robust prediction from external multi-source data as a prior, towards which we shrink an algorithm based on relatively few target data, proves to be a coherent and promising approach.
Finally, our proposed framework for evaluating external training data utility is highly general and we would welcome its application to other (multi-source) domain adaptation learning problems.
@@
Our study uses harmonized data and resources expanding the ricu R-package (Bennett et al., 2023) from prior work by Burger et al. (2024); Burger et al. (2025), who describe a large-scale multi-center ICU dataset. This extended version of ricu includes additional code and configuration files to harmonize data from additional data sources (NWICU, PICdb, and Zigong) and to extract additional variables and will be published together with Burger et al. (2025). See https://github.com/ratschlab/icarefm for details.
@@
Funding
Malte Londschien is partially supported by the ETH Foundations of Data Science.
Conflicts of interest
The authors declare no conflicts of interest.
@@ Acknowledgements
We gratefully acknowledge early access to the harmonized dataset and ricu resources provided by Burger et al. (2024); Burger et al. (2025). We would like to thank Cyrill Scheidegger, Dinara Veshchezerova, Dominik Rothenhäusler, Gianna Wolfisberg, Hugo Yèche, Jonas Peters, Julia Kostin, Markus Ulmer, Maybritt Schillinger, Michael Law, Nicolas Bennet, Niklas Pfister, Olga Mineeva, Patrick Rockenschaub, Paola Malsot, Sorawit Saengkyongam, and Yuansi Chen for helpful discussions and comments. Finally, we thank the editor, associate editor and reviewers for their constructive comments.
@@ A.1 Tasks
Acute kidney injury is defined as AKI 3 according to the KDIGO guidelines (Acute Kidney Injury Work Group, 2012). These are: (i) A patient has an increase in creatinine of a factor of 3 relative to their 7 day baseline; (ii) A patient has acute kidney injury level 1 according to the KDIGO guidelines and has a creatinine value of at least 4.0 mmol/L; (iii) A patient has anuria, that is, has not produced urine over the last 12 hours; (iv) A patient had an average relative urine rate below 0.3ml/kg/h over the last 24 hours; (v) The patient has started renal replacement therapy. We say that a patient is experiencing an acute kidney injury event if at least one of the conditions (i) - (v) apply. As urine rate is only measured sparsely for NWICU, PICdb, and Zigong, we define a negative kidney injury event as (i, ii) creatinine is measured and low, (iii, iv) urine rate is measured and normal, or not measured, (v) the patient is not receiving renal replacement therapy. Again, requiring that urine rate is measured and normal would result in very few negative events for the datasets where urine rate is only very sparsely measured.
@@ A.2 Feature engineering
We use an 8 hour horizon for feature engineering for the prediction tasks of log(lactate) and circulatory failure and a 24 hour horizon for the feature engineering for the prediction tasks of log(creatinine), acute kidney injury.
@@ B.1 Details on the probit anchor loss
Recall that $\Phi$ and $\varphi$ are the Gaussian distribution’s cumulative distribution function and probability density function. For scores $f\in\mathbb{R}^{n}$ and outcome $y\in\{-1,1\}^{n}$, the negative log-likelihood is $\sum_{i=1}^{n}\log(\Phi(y_{i}f_{i}))$ with gradient $r(f):=-y\cdot\varphi(f)/\Phi(yf)$. We use the gradient as (score) residuals (Kook et al., 2022). The probit anchor loss with parameter $\gamma$ is
@@
$H(f,y)=\operatorname{diag}(\dot{r}(f)+(\gamma-1)P_{A}r\cdot\ddot{r}(f))+(\gamma-1)\cdot\operatorname{diag}(\dot{r}(f))P_{A}\operatorname{diag}(\dot{r}(f)).$ (7)
@@ B.2 The logistic anchor loss is not convex
Write $\sigma(f)=1/(1+e^{-f})$. For scores $f\in\mathbb{R}^{n}$ and outcomes $y\in\{-1,1\}^{n}$, the negative log-likelihood using the logistic link $\sigma$ is $\sum_{i=1}^{n}\log(\sigma(y_{i}f_{i}))$ with gradient $r(f):=y\cdot\sigma(-yf)$. Using the gradient as (score) residuals (Kook et al., 2022), the logistic anchor loss with parameter $\gamma$ is
@@ Appendix C Additional figures
Figure 28: Regime transitions for linear models as described in section˜3.6 and figure˜2. denotes the regime transition i $\to$ $i$,   the regime transition $i$ $\to$ iii, and denotes the external data’s value.